OpenStudy (anonymous):

need help! cosx-sinex =1 for 0<= x<=360 i got answer like x= 0 ,180,270 but only 0 and 270 are taken as roots of this equation . why 180 is discarded ?

OpenStudy (ash2326):

\[\cos x-\sin x=1\] put x=180 \[\cos 180-\sin 180=\] \[-1-0=-1\ne 1\]

OpenStudy (ash2326):

do you get this part?

OpenStudy (anonymous):

yep !

OpenStudy (ash2326):

so 180 is not an answer, right?

OpenStudy (anonymous):

yah! now i got it !

OpenStudy (anonymous):

thanks !

OpenStudy (ash2326):

ok =)